# Homomorphism counting for immersion-closed classes is not isomorphism

### v1

@@ 4 A tight bound for oddomorphisms and treewidth
To prove Theorem 1.10, suppose towards a contradiction that $\operatorname{tw}(G)\leq t$, $G$ has a $(t+2)$-oddomorphism $f$, and subject to these conditions, $G$ minimizes $|V(G)|+|E(G)|$. As removing edges does not increase the treewidth, Lemma 2.5 implies that for any two colour classes $C_{i},C_{j}$, we have that $G[C_{i},C_{j}]$ is acyclic. Let $(T,\{\beta_{t}\}_{t\in V(T)})$ be a tree decomposition of $G$ of width at most $t$. Fix a vertex $r\in V(T)$ and view $T$ as a tree rooted at $r$.
For every $f$-odd vertex $u$ in $G$, let $D(u,r)$ be the shortest distance in $T$ from $r$ to a bag containing $u$. Let $w$ be an $f$-odd vertex which maximizes $D(u,r)$ over all $f$-odd vertices $u$ (picking arbitrarily if there are more than one). Let $\beta_{v}$ be the bag which attains $D(w,r)$. Note that $v\neq r$ because there are at least $t+2$ ff-odd vertices and $|\beta_{r}|\leq t+1$.
As $w$ is $f$-odd, for each of the $t+1$ colours distinct from $f(w)$, $w$ has a neighbour of that colour. Since $w\in\beta_{v}$ and $|\beta_{v}|\leq t+1$, there exists a colour $c^{\prime}$ such that no vertex in $\beta_{v}$ has colour $c^{\prime}$. Let $w^{\prime}$ be a neighbour of $w$ of colour $c^{\prime}$. Note that if $w^{\prime}$ appears in a bag of the component of TvT-v containing $r$, then $w$ also appears in that bag by the definition of tree decompositions, contradicting our choice of $\beta_{v}$. Hence, $w^{\prime}$ appears in a bag of a component of TvT-v not containing $r$. Let $c=f(w)$.
Observe that in $G[C_{c},C_{c^{\prime}}]$ there is a path starting at $w$, going through $w^{\prime}$ and ending in some $f$-odd vertex $w^{\prime\prime}$. Let $P$ be such a path. If PwP-w intersects $\beta_{v}$, then by our choice of the colour $c^{\prime}$, there is a vertex $q\in\beta_{v}$ such that $q\neq w$ and $q$ has colour $c$. In this case, consider the merger of $q$ and $w$ and let $G^{\prime}$ be the resulting graph. Then $|V(G^{\prime})|<|V(G)|$ and, by Lemma 2.4, $G^{\prime}$ admits a $(t+2)$-oddomorphism. For each $t\in V(T)$, let $\beta_{t}^{\prime}$ be the bag obtained from $\beta_{t}$ by replacing $q$ and $w$ (if they appear) with the merged vertex. Since $q$ and $w$ are in a common bag of $(T,\{\beta_{t}\}_{t\in V(T)})$, we have that $(T,\{\beta_{t}^{\prime}\}_{t\in V(T)})$ is a tree decomposition of $G^{\prime}$ of width $t$, contradicting the minimality of the counterexample $G$.
@@ 5 Homomorphism counting and oddomorphisms
For any $n\in\mathbb{N}$, there exists non-isomorphic graphs $H$ and $H^{\prime}$ such that $\hom(F,H)\neq\hom(F,H^{\prime})$ implies that $F$ contains a $K_{n}$-immersion.
@@
Now our goal is to prove Theorem 5.1.
We need to utilize a theorem from [22], and as such we need to extend the notion of oddomorphism from cliques to all graphs.
@@
Let $f:V(G)\rightarrow H$ be a homomorphism. A vertex $u\in V(G)$ is $f$-odd (respectively, $f$-even) if $N_{G}(v)$ contains an odd (respectively, even) number of vertices of each colour in $S\setminus\{f(v)\}$for every $v\in V(H)\cap N_{H}(f(u))$, the number of vertices $w$ in $N_{G}(u)$ such that $f(w)=v$ is odd. Similarly, $u$ is $f$-even if for every $v\in V(H)\cap N(f(u))$, the number of vertices $w$ in $N_{G}(u)$ such that $f(w)=v$ is even.
@@
Let $f:G\to F$. We say $f$ is a weak oddomorphism if there is a (not necessarily induced) subgraph $G^{\prime}$ of $F$ such that the restriction of $f$ to $G^{\prime}$ is an oddomorphism from $G^{\prime}$ to $F$.
@@
Let $G$ be a connected graph. Given $G$, there are non-isomorphic graphs $G_{0}$ and $G_{1}$ constructed from $G$ such that $\hom(F,G_{0})\geq\hom(F,G_{1})$ with strict inequality if and only if there exists a weak oddomorphism from $F$ to $G$. Moreover, if such a weak oddomorphism $\phi$ does exist, then there is a connected subgraph $F^{\prime}$ of $F$ such that $\phi|_{V(F^{\prime})}$ is an oddomorphism from $F^{\prime}$ to $G$.
@@
Let $n\in\mathbb{N}$ be given. Let $G=K_{{n\choose 2}(7n+7)}$. By Theorem 5.6, there are non-isomorphic graphs $G_{0},G_{1}$ such that $\hom(F,G_{0})\geq\hom(F,G_{1})$ and strict inequality holds if and only if $F$ has a weak oddomorphism to $G$. As $G_{1}$ and $G_{0}$ are not isomorphic, by Theorem 1.1 there is a graph $F$ such that $\hom(F,G_{0})\neq\hom(F,G_{1})$. For all such graphs $F$, we have that $F$ admits a weak oddomorphism to $G$. Thus there is a subgraph $F^{\prime}$ of $F$ such that $F^{\prime}\xrightarrow{\text{odd}}G$. But by Theorem 1.8, this implies $F^{\prime}$ has a $K_{t}$-immersion, and hence $F$ has a $K_{t}$-immersion, thus implying Theorem 5.1. ∎
We note that we could perform a similar proof to show that if $\mathcal{M}$ is the class of graphs with treewidth at most $k$, then $_{\mathcal{M}}$ is not isomorphism, however this is already known [20].

### v2

@@ 4 A tight bound for oddomorphisms and treewidth
In this section we prove Theorem 1.10.
Suppose towards a contradiction that $\operatorname{tw}(G)\leq t-2$, $G$ has a $t$-oddomorphism $f$, and subject to these conditions, $G$ minimizes $|V(G)|+|E(G)|$. As removing edges does not increase the treewidth, Lemma 2.5 implies that for any two colour classes $C_{i},C_{j}$, we have that $G[C_{i},C_{j}]$ is acyclic. Let $(T,\{\beta_{x}\}_{x\in V(T)})$ be a tree decomposition of $G$ of width at most t2t-2. Fix a vertex $r\in V(T)$ and view $T$ as a tree rooted at $r$.
For every $f$-odd vertex $u$ in $G$, let $D(u,r)$ be the shortest distance in $T$ from $r$ to a bag containing $u$. Let $w$ be an $f$-odd vertex which maximizes $D(u,r)$ over all $f$-odd vertices $u$ (picking arbitrarily if there are more than one). Let $\beta_{v}$ be the bag which attains $D(w,r)$. Note that $v\neq r$ because there are at least $t$ $f$-odd vertices and $|\beta_{r}|\leq t-1$.
As $w$ is $f$-odd, for each of the t1t-1 colours distinct from $f(w)$, $w$ has a neighbour of that colour. Since $w\in\beta_{v}$ and $|\beta_{v}|\leq t-1$, there exists a colour $c^{\prime}$ such that no vertex in $\beta_{v}$ has colour $c^{\prime}$. Let $w^{\prime}$ be a neighbour of $w$ of colour $c^{\prime}$. Note that if $w^{\prime}$ appears in a bag of the component of TvT-v containing $r$, then $w$ also appears in that bag by the definition of tree decompositions, contradicting our choice of $\beta_{v}$. Hence, $w^{\prime}$ appears in a bag of a component of TvT-v not containing $r$. Let $c=f(w)$.
Observe that in $G[C_{c},C_{c^{\prime}}]$ there is a path starting at $w$, going through $w^{\prime}$ and ending in some $f$-odd vertex $w^{\prime\prime}$. Let $P$ be such a path. If PwP-w intersects $\beta_{v}$, then by our choice of the colour $c^{\prime}$, there is a vertex $q\in\beta_{v}$ such that $q\neq w$ and $q$ has colour $c$. In this case, consider the merger of $q$ and $w$ and let $G^{\prime}$ be the resulting graph. Then $|V(G^{\prime})|<|V(G)|$ and, by Lemma 2.4, $G^{\prime}$ admits a $t$-oddomorphism. For each $x\in V(T)$, let $\beta_{x}^{\prime}$ be the bag obtained from $\beta_{x}$ by replacing $q$ and $w$ (if they appear) with the merged vertex. Since $q$ and $w$ are in a common bag of $(T,\{\beta_{x}\}_{x\in V(T)})$, we have that $(T,\{\beta_{x}^{\prime}\}_{x\in V(T)})$ is a tree decomposition of $G^{\prime}$ of width t2t-2, contradicting the minimality of the counterexample $G$.
@@ 5 Homomorphism counting and oddomorphisms
For any $n\in\mathbb{N}$, there exists non-isomorphic graphs $H$ and $H^{\prime}$ such that for every graph $F$, $\hom(F,H)\neq\hom(F,H^{\prime})$ implies that $F$ contains a $K_{n}$-immersion.
@@
Hence, our goal is to prove Theorem 5.1. We first extend the notion of oddomorphism from cliques to all graphs in order to utilize a theorem from [23].
@@
Let $f:V(G)\rightarrow V(H)$ be a homomorphism. A vertex $u\in V(G)$ is $f$-odd (respectively, $f$-even) if for every $v\in V(H)\cap N_{H}(f(u))$, the number of vertices $w$ in $N_{G}(u)$ such that $f(w)=v$ is odd (respectively, even)
@@
Let $F$ and $G$ be graphs and let $f$ be a homomorphism from $F$ to $G$. We say that $f$ is a weak oddomorphism from $F$ to $G$ if there is a (not necessarily induced) subgraph $F^{\prime}$ of $F$ such that the restriction of $f$ to $V(F^{\prime})$ is an oddomorphism from $F^{\prime}$ to $G$.
@@
Let $G$ be a connected graph. Then there are non-isomorphic graphs $G_{0}$ and $G_{1}$ such that $\hom(F,G_{0})\geq\hom(F,G_{1})$ for every graph $F$ with strict inequality if and only if there exists a weak oddomorphism from $F$ to $G$. Moreover, if such a weak oddomorphism $f$ exists, then there is a connected subgraph $F^{\prime}$ of $F$ such that $f|_{V(F^{\prime})}$ is an oddomorphism from $F^{\prime}$ to $G$.
@@
Let $n\in\mathbb{N}$ be given. Let $G=K_{{n\choose 2}(7n+7)}$. By Theorem 5.6, there are non-isomorphic graphs $G_{0},G_{1}$ such that $\hom(F,G_{0})\geq\hom(F,G_{1})$ for every graph $F$ with strict inequality if and only if $F$ has a weak oddomorphism to $G$. As $G_{0}$ and $G_{1}$ are not isomorphic, by Theorem 1.1 there is a graph $F$ such that $\hom(F,G_{0})\neq\hom(F,G_{1})$. For all such graphs $F$, we have that $F$ admits a weak oddomorphism to $G$. Thus there is a subgraph $F^{\prime}$ of $F$ such that $F^{\prime}\xrightarrow{\text{odd}}G$. But by Theorem 1.8, this implies $F^{\prime}$ has a $K_{t}$-immersion, and hence $F$ has a $K_{t}$-immersion, thus implying Theorem 5.1. ∎
A similar proof using Theorem 1.10 shows that if $\mathcal{M}$ is the class of graphs with treewidth at most $k$, then $_{\mathcal{M}}$ is not isomorphism. Since this approach was already taken in [21] and this is a known result, we omit this proof.